OpenStudy (anonymous):

Solve the formula P = 2l + 2w for w

OpenStudy (anonymous):

P = 2l + 2w i.e. P - 2I = 2w \[\huge w= \frac{P-2I}{2}\] \[\rightarrow \huge w= \frac{P}{2} - \frac{2I}{2} \] \[\rightarrow \huge w= \frac{P}{2} - I\] @kpowell032

OpenStudy (anonymous):

you have to find L first.